# Extremal Process of Last Progeny Modified Branching Random Walks

### v1

@@ 3.1 BRW below the boundary case
$\left|g_{\varphi}(x)-c_{\varphi}(\theta)Le^{\theta x}\right|\leq\varepsilon e^{\theta x}.$ (3.1)
Observe that for all $a,b>0$, we have $|e^{-a}-e^{-b}|\leq|a-b|\wedge 1$ using that $\exp$ is $1$-Lipschitz on $\mathbb{R}_{-}$. Therefore,
$\operatorname{\mathbf{E}}\left(\left|e^{-{\langle\tau_{-m_{n}}\mathcal{X}_{n},g_{\varphi}\rangle}}-e^{-c_{\varphi}(\theta)W_{n}(\theta)}\right|\right)$ $\leq\operatorname{\mathbf{E}}\left(\left|{\langle\tau_{-m_{n}}\mathcal{X}_{n},g_{\varphi}\rangle}-c_{\varphi}(\theta)W_{n}(\theta)\right|\wedge 1\right)$
$\leq\operatorname{\mathbf{E}}\left(\left|\sum_{|u|=n}h_{\varphi}(S_{u}-m_{n})\right|\wedge 1\right),$
where $h_{\varphi}(x)=g_{\varphi}(x)-c_{\varphi}(\theta)Le^{\theta x}$, using that $e^{-\theta m_{n}}=L^{-1}e^{-n\kappa(\theta)}$. By (3.1), we have
$\operatorname{\mathbf{E}}\left(\mathbbm{1}_{\left\{M_{n}\leq m_{n}-A\right\}}\left|\sum_{|u|=n}h_{\varphi}(S_{u}-m_{n})\right|\wedge 1\right)\leq\varepsilon L^{-1}\operatorname{\mathbf{E}}(W_{n}(\theta)),$
@@
$\operatorname{\mathbf{E}}\left(\left|e^{-{\langle\tau_{-m_{n}}\mathcal{X}_{n},g_{\varphi}\rangle}}-e^{-c_{\varphi}(\theta)W_{n}(\theta)}\right|\right)\leq\varepsilon L^{-1}+\operatorname{\mathbf{P}}(M_{n}\geq m_{n}-A).$
Letting $n\to\infty$, then $\varepsilon\to 0$, we conclude, by Lebesgue’s dominated convergence theorem and (2.2), that
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(e^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right)$ $=\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(e^{-c_{\varphi}(\theta)\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}}\right)$
$=\operatorname{\mathbf{E}}\left(\exp\left(-W_{\infty}(\theta)\int\theta e^{-\theta z}(1-e^{-\varphi(z)})\,\mathrm{d}z\right)\right).$
To complete the proof, it is then enough to observe that this limit is the Laplace transform of the PPP($\theta W_{\infty}(\theta)e^{-\theta z}\,\mathrm{d}z$). ∎
@@
$\operatorname{\mathbf{E}}\left(\exp\left(-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}\right)\right)=\operatorname{\mathbf{E}}\left(\exp\left(-\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\right)\right),$
@@
$(1-\tfrac{2}{c_{1}^{1/\alpha}}\eta)^{|\alpha|}c_{1}c_{\varphi}(\theta)\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|\leq n\eta\right\}}$
@@
$\qquad\qquad\leq(1-\tfrac{2}{c_{1}^{1/\alpha}}\eta)^{-|\alpha|}c_{1}c_{\varphi}(\theta)\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|\leq n\eta\right\}}.$
Here we use the Uniform Convergence Theorem [9, Theorem 1.5.2], which states that $\lim_{z\to\infty}\frac{L(\lambda z)}{L(z)}=\lambda^{\alpha}$ holds uniformly on compact subsets of $(0,\infty)$. Note that there exists $r>0$ so that
$[(1-\tfrac{3}{c_{1}^{1/\alpha}}\eta)^{|\alpha|},(1-\tfrac{3}{c_{1}^{1/\alpha}}\eta)^{-|\alpha|}]\subset[1-r\eta,1+r\eta].$
Using Lemma 2.5, together with the law of large numbers, we observe that
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|>n\eta\right\}}\right)=\lim_{n\to\infty}\operatorname{\mathbf{P}}(|T_{n}-\operatorname{\mathbf{E}}(T_{n})|>n\eta)=0.$ (3.2)
@@
$\lim_{n\to\infty}\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|\leq n\eta\right\}}=W_{\infty}(\theta)\quad\text{ in probability.}$.
@@
converges to 00 as $n\to\infty$. In order to show that $\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})$ converges to $c_{1}c_{\varphi}(\theta)W_{\infty}(\theta)$ in probability and complete the proof of Theorem 1.3, we now show that
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|>n\eta\right\}}=0\quad\text{ in probability.}$. (3.3)
On the one hand, using that $\theta\mapsto\kappa(\theta)$ is $\mathcal{C}^{2}$ on $(\theta-\delta,\theta+\delta)$, and that $\theta\kappa^{\prime}(\theta)-\kappa(\theta)<0$, we observe that $\vartheta\mapsto\kappa(\vartheta)/\vartheta$ is decreasing on $[\theta,\theta+\delta]$ for $\delta$ small enough. Thus, by Lemma 2.1, there exists $\varepsilon>0$ such that
[…]
@@
$\qquad\leq 2c_{\varphi}(\theta)\sum_{|u|=n}L(m_{n}-S_{u})\frac{e^{\theta S_{u}-n\kappa(\theta)}}{L(n)}\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta,S_{u}-m_{n}<-n\varepsilon/2\right\}}.$
Therefore, using again [9, Theorem 1.5.2], there exists a constant $C$ depending on $\varepsilon$ and $\eta$ such that almost surely, for all $n$ large enough
$\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta\right\}}\leq C\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta\right\}}.$
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta\right\}}=0\quad\text{ in probability.}$. (3.4)
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)<-n\eta\right\}}\mathbbm{1}_{\left\{S_{u}-m_{n}>-nB\right\}}=0\quad\text{ in probability.}$. (3.5)
@@
$g_{\varphi}(-x)\leq 2c_{\varphi}(\theta)L(x)e^{-\theta x}\leq 4c_{\varphi}(\theta)e^{-(\theta-\delta/2)x}.$
@@
$\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}<-nB\right\}}$ $\leq 4c_{\varphi}(\theta)\sum_{|u|=n}e^{(\theta-\delta/2)(S_{u}-m_{n})}\mathbbm{1}_{\left\{S_{u}-m_{n}<-nB\right\}}$
$\leq 4c_{\varphi}(\theta)W_{n}(\theta-\delta)e^{n\kappa(\theta-\delta)-(\theta-\delta)m_{n}-\delta nB/2}.$
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}<-nB\right\}}=0\quad\text{ in probability.}$. (3.6)
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})=c_{1}c_{\varphi}(\theta)W_{\infty}(\theta)\quad\text{ in probability.}$.
Using the dominated convergence theorem together with (2.2), we conclude that for all for all $\varphi\in\mathcal{T}$,
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(e^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right)=\operatorname{\mathbf{E}}\left(\exp\left(-c_{1}W_{\infty}(\theta)c_{\varphi}(\theta)\right)\right).$
We can now complete the proof, observing that the right-hand side is the Laplace transform of the PPP($c_{1}\theta W_{\infty}(\theta)e^{-\theta x}\,\mathrm{d}x$). ∎
3.2 BRW in the boundary case
@@
$\mathcal{G}_{n}(A)=\{|u|=n:S_{u_{k}}\leq k\kappa^{\prime}(\theta)+A,k\leq n\},$
the set of particles that stayed at all times below the line $x\mapsto x\kappa^{\prime}(\theta)+A$. Using that $\lim_{n\to\infty}M_{n}-n\kappa^{\prime}(\theta)=-\infty$ a.s. we observe that almost surely, for $A$ large enough, we have $\mathcal{G}_{n}(A)=\{|u|=n\}$. In other words, writing
@@
we observe that on the event $\mathcal{S}_{A}$, we have $\mathcal{G}_{n}(A)=\{|u|=n\}$ for all $n\in\mathbb{N}$, and that $\lim_{A\to\infty}\operatorname{\mathbf{P}}(\mathcal{S}_{A})=1$.
As a first step towards the proof of Theorem 1.4, we show that no particle in $\mathcal{G}_{n}(A)$ above position $m_{n}-\varepsilon n^{1/2}$ contribute to the extremal process with high probability, as soon as $\varepsilon>0$ is small enough.
@@
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}(\exists|u|=n:S_{u}\geq m_{n}-\varepsilon n^{1/2},S_{u}+Y_{u}\geq m_{n}-B)=0.$
@@
We first recall from [18, Theorem 1.1], that
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}(M_{n}\geq n\kappa^{\prime}(\theta)-\frac{3}{2\theta}\log n+\varepsilon^{-1})=0.$
@@
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}(M_{n}\geq m_{n}-a_{n}+\varepsilon^{-1})=0.$
@@
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}\left(\exists u\in\mathcal{G}_{n}(A):\begin{array}[]{l}m_{n}-S_{u}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],\\ S_{u}+Y_{u}\geq m_{n}-B\end{array}\right)=0.$ (3.7)
@@
[…]
@@
$\operatorname{\mathbf{E}}\left(\sum_{u\in\mathcal{G}_{n}(A)}\mathbbm{1}_{\left\{m_{n}-S_{u}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}]\right\}}L(m_{n}-S_{u})e^{\theta(S_{u}-m_{n})}\right)$
[…]
where $\widehat{T}_{k}=T_{k}-k\kappa^{\prime}(\theta)$ and $\widehat{m}_{n}=m_{n}-n\kappa^{\prime}(\theta)$. Let $\rho\in(0,\alpha+2)$, we define $\mathfrak{L}:x\mapsto x^{\rho-\alpha}L(x)$. Observe that $\mathfrak{L}$ is a regularly varying function at $\infty$ with index $\rho>0$. Therefore, by [9, Theorem 1.5.2], for all $\delta>0$, there exists $N_{\delta}$ such that for all $x>N_{\delta}$,
@@
$\operatorname{\mathbf{E}}\left(\frac{L(\widehat{m}_{n}-\widehat{T}_{n})}{L(n^{1/2})}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right)\\ \leq\operatorname{\mathbf{E}}\left(\left(\frac{\widehat{m}_{n}-\widehat{T}_{n}}{n^{1/2}}\right)^{\alpha}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right)\\ +\delta\operatorname{\mathbf{E}}\left(\left(\frac{\widehat{m}_{n}-\widehat{T}_{n}}{n^{1/2}}\right)^{\alpha-\rho}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right).$
@@
Using e.g. [3, Lemma 4.1], there exist $C>0$ and $h>0$ such that for all $n\in\mathbb{N}$, $a\geq 0$ and $b\geq-a$, we have
$\operatorname{\mathbf{P}}(-\widehat{T}_{n}\in[b,b+h],-\widehat{T}_{j}\geq-a,j\leq n)\leq C\frac{((a+1)\wedge n^{1/2})((a+b+1)\wedge n^{1/2})}{n^{3/2}}.$
@@
$\operatorname{\mathbf{E}}\left(\left(\widehat{m}_{n}-\widehat{T}_{n}\right)^{\gamma}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right)$
[…]

### v2

@@ 3.1. BRW below the boundary case
$\left|g_{\varphi}(x)-c_{\varphi}(\theta)L\mathrm{e}^{\theta x}\right|\leq\varepsilon\mathrm{e}^{\theta x}.$ (3.1)
Observe that for all $a,b\geq 0$, we have $|\mathrm{e}^{-a}-\mathrm{e}^{-b}|\leq|a-b|\wedge 1$ using that $\exp$ is $1$-Lipschitz on $\mathbb{R}_{-}$. Therefore,
$\operatorname{\mathbf{E}}\left[\left|\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{X}_{n},g_{\varphi}\rangle}}-\mathrm{e}^{-c_{\varphi}(\theta)W_{n}(\theta)}\right|\right]$ $\leq\operatorname{\mathbf{E}}\left[\left|{\langle\tau_{-m_{n}}\mathcal{X}_{n},g_{\varphi}\rangle}-c_{\varphi}(\theta)W_{n}(\theta)\right|\wedge 1\right]$
$\leq\operatorname{\mathbf{E}}\left[\left|\sum_{|u|=n}h_{\varphi}(S_{u}-m_{n})\right|\wedge 1\right],$
where $h_{\varphi}(x)=g_{\varphi}(x)-c_{\varphi}(\theta)L\mathrm{e}^{\theta x}$, using that $\mathrm{e}^{-\theta m_{n}}=L^{-1}\mathrm{e}^{-n\kappa(\theta)}$. By (3.1), we have
$\operatorname{\mathbf{E}}\left[\mathbbm{1}_{\left\{M_{n}\leq m_{n}-A\right\}}\left|\sum_{|u|=n}h_{\varphi}(S_{u}-m_{n})\right|\wedge 1\right]\leq\varepsilon L^{-1}\operatorname{\mathbf{E}}[W_{n}(\theta)],$
@@
$\operatorname{\mathbf{E}}\left[\left|\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{X}_{n},g_{\varphi}\rangle}}-\mathrm{e}^{-c_{\varphi}(\theta)W_{n}(\theta)}\right|\right]\leq\varepsilon L^{-1}+\operatorname{\mathbf{P}}(M_{n}\geq m_{n}-A).$
Letting $n\to\infty$ and then $\varepsilon\downarrow 0$, we conclude, by first using Lemma 2.1 and then using Lebesgue’s dominated convergence theorem and (2.2), that
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right]$ $=\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-c_{\varphi}(\theta)\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}}\right]$
$=\operatorname{\mathbf{E}}\left[\exp\left(-W_{\infty}(\theta)\int\theta\mathrm{e}^{-\theta z}(1-\mathrm{e}^{-\varphi(z)})\,\mathrm{d}z\right)\right].$
[…]
@@
$\operatorname{\mathbf{E}}\left[\exp\left(-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}\right)\right]=\operatorname{\mathbf{E}}\left[\exp\left(-\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\right)\right],$
@@
$(1-\tfrac{2}{c_{1}^{1/\alpha}}\eta)^{|\alpha|}c_{1}c_{\varphi}(\theta)\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|\leq n\eta\right\}}$
@@
$\qquad\qquad\leq(1-\tfrac{2}{c_{1}^{1/\alpha}}\eta)^{-|\alpha|}c_{1}c_{\varphi}(\theta)\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|\leq n\eta\right\}}.$
Here we use the Uniform Convergence Theorem (Bingham et al., 1987, Theorem 1.5.2), which states that $\lim_{z\to\infty}\frac{L(\lambda z)}{L(z)}=\lambda^{\alpha}$ holds uniformly on compact subsets of $(0,\infty)$. Note that there exists $r>0$ so that
$(1-\tfrac{3}{c_{1}^{1/\alpha}}\eta)^{|\alpha|},(1-\tfrac{3}{c_{1}^{1/\alpha}}\eta)^{-|\alpha|}]\subset[1-r\eta,1+r\eta].$
Using Lemma 2.4, together with the law of large numbers for the random walk $(T_{n},n\geq 0)$, we observe that
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|>n\eta\right\}}\right]=\lim_{n\to\infty}\operatorname{\mathbf{P}}\left(\left|T_{n}-\operatorname{\mathbf{E}}[T_{n}]\right|>n\eta\right)=0.$ (3.2)
@@
$\lim_{n\to\infty}\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|\leq n\eta\right\}}=W_{\infty}(\theta)\quad\text{ in probability.}$
@@
converges to 0 as $n\to\infty$. In order to show that $\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})$ converges to $c_{1}c_{\varphi}(\theta)W_{\infty}(\theta)$ in probability and complete the proof of Theorem 1.3, we now show that
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{|S_{u}-n\kappa^{\prime}(\theta)|>n\eta\right\}}=0\quad\text{ in probability.}$ (3.3)
On the one hand, using that $\vartheta\mapsto\kappa(\vartheta)$ is $\mathcal{C}^{2}$ on $(\theta-\delta,\theta+\delta)$, and that $\theta\kappa^{\prime}(\theta)-\kappa(\theta)<0$, we observe that $\kappa(\vartheta)/\vartheta$ is decreasing on $\theta,\theta+\delta]$ for sufficiently small $\delta$, which implies $\frac{\kappa(\theta)}{\theta}>\frac{\kappa(\theta+\delta)}{\theta+\delta}$. Hence, by Lemma 2.1, there exists $\varepsilon>0$ such that
$\lim_{n\to\infty}M_{n}-n\left(\frac{\kappa(\theta)}{\theta}-\varepsilon\right)=-\infty\quad\text{a.s.}$
@@
$\qquad\leq 2c_{\varphi}(\theta)\sum_{|u|=n}L(m_{n}-S_{u})\frac{\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}}{L(n)}\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta,S_{u}-m_{n}<-n\varepsilon/2\right\}}.$
Therefore, using again the Uniform Convergence Theorem (Bingham et al., 1987, Theorem 1.5.2), there exists a constant $C$ depending on $\varepsilon$ and $\eta$ such that almost surely, for all $n$ large enough
$\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta\right\}}\leq C\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta\right\}}.$
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)>n\eta\right\}}=0\quad\text{ in probability.}$ (3.4)
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-n\kappa^{\prime}(\theta)<-n\eta\right\}}\mathbbm{1}_{\left\{S_{u}-m_{n}>-nB\right\}}=0\quad\text{ in probability.}$ (3.5)
@@
$g_{\varphi}(-x)\leq 2c_{\varphi}(\theta)L(x)\mathrm{e}^{-\theta x}\leq 4c_{\varphi}(\theta)\mathrm{e}^{-(\theta-\delta/2)x}.$
@@
$\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}<-nB\right\}}$ $\leq 4c_{\varphi}(\theta)\sum_{|u|=n}\mathrm{e}^{(\theta-\delta/2)(S_{u}-m_{n})}\mathbbm{1}_{\left\{S_{u}-m_{n}<-nB\right\}}$
$\leq 4c_{\varphi}(\theta)W_{n}(\theta-\delta)\mathrm{e}^{n\kappa(\theta-\delta)-(\theta-\delta)m_{n}-\delta nB/2}.$
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}<-nB\right\}}=0\quad\text{ in probability.}$ (3.6)
@@
$\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})=c_{1}c_{\varphi}(\theta)W_{\infty}(\theta)\quad\text{ in probability.}$
Using the dominated convergence theorem together with (2.2), we conclude that for all $\varphi\in\mathcal{T}$,
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right]=\operatorname{\mathbf{E}}\left[\exp\left(-c_{1}W_{\infty}(\theta)c_{\varphi}(\theta)\right)\right].$
[…]
3.2. BRW in the boundary case
@@
$\mathcal{G}_{n}(A)=\{u\in\mathcal{U}:|u|=n,\;S_{u_{k}}\leq k\kappa^{\prime}(\theta)+A\quad\forall\,k\leq n\},$
the set of particles that stayed at all times below the line $x\mapsto x\kappa^{\prime}(\theta)+A$. Using that $\lim_{n\to\infty}M_{n}-n\kappa^{\prime}(\theta)=$-$\infty$ a.s. we observe that almost surely, for $A$ large enough, we have $\mathcal{G}_{n}(A)=\{u\in\mathcal{U}:|u|=n\}$. In other words, writing
@@
we observe that on the event $\mathcal{S}_{A}$, we have $\mathcal{G}_{n}(A)=\{u\in\mathcal{U}:|u|=n\}$ for all $n\in\mathbb{N}$, and that $\lim_{A\to\infty}\operatorname{\mathbf{P}}(\mathcal{S}_{A})=1$.
As a first step towards the proof of Theorem 1.4, we show that no particle in $\mathcal{G}_{n}(A)$ above position $m_{n}-\varepsilon n^{1/2}$ contributes to the extremal process with high probability, as soon as $\varepsilon>0$ is small enough.
@@
$\lim_{\varepsilon\downarrow 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}(\exists u\in\mathcal{U}:|u|=n,S_{u}\geq m_{n}-\varepsilon n^{1/2},S_{u}+Y_{u}\geq m_{n}-B)=0.$
@@
We first recall from (Mallein, 2016, Theorem 1.1), that
$\lim_{\varepsilon\downarrow 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}(M_{n}\geq n\kappa^{\prime}(\theta)-\frac{3}{2\theta}\log n+\varepsilon^{-1})=0.$
@@
$\lim_{\varepsilon\downarrow 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}(M_{n}\geq m_{n}-a_{n}+\varepsilon^{-1})=0.$
@@
$\lim_{\varepsilon\downarrow 0}\limsup_{n\to\infty}\operatorname{\mathbf{P}}\left(\exists u\in\mathcal{G}_{n}(A):\begin{array}[]{l}m_{n}-S_{u}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],\\ S_{u}+Y_{u}\geq m_{n}-B\end{array}\right)=0.$ (3.7)
@@
$\qquad\leq\operatorname{\mathbf{E}}\left[\sum_{u\in\mathcal{G}_{n}(A)}\mathbbm{1}_{\left\{m_{n}-S_{u}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}]\right\}}\nu([m_{n}-B-S_{u},\infty))\right]$
$\qquad\leq 2\operatorname{\mathbf{E}}\left[\sum_{u\in\mathcal{G}_{n}(A)}\mathbbm{1}_{\left\{m_{n}-S_{u}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}]\right\}}L(m_{n}-S_{u})\mathrm{e}^{\theta(S_{u}+B-m_{n})}\right]$
@@
$\operatorname{\mathbf{E}}\left[\sum_{u\in\mathcal{G}_{n}(A)}\mathbbm{1}_{\left\{m_{n}-S_{u}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}]\right\}}L(m_{n}-S_{u})\mathrm{e}^{\theta(S_{u}-m_{n})}\right]$
$\qquad=n^{1/2}\operatorname{\mathbf{E}}\left[\sum_{u\in\mathcal{G}_{n}(A)}\frac{L(m_{n}-S_{u})}{L(n^{1/2})}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{m_{n}-S_{u}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}]\right\}}\right]$
$\qquad=n^{1/2}\operatorname{\mathbf{E}}\left[\frac{L(\widehat{m}_{n}-\widehat{T}_{n})}{L(n^{1/2})}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right],$
where $\widehat{T}_{k}=T_{k}-k\kappa^{\prime}(\theta)$ and $\widehat{m}_{n}=m_{n}-n\kappa^{\prime}(\theta)$. Let $\rho\in(0,\alpha+2)$, we define $\mathfrak{L}:x\mapsto x^{\rho-\alpha}L(x)$. Observe that $\mathfrak{L}$ is a regularly varying function at $\infty$ with index $\rho>0$. Therefore, by the Uniform Convergence Theorem (Bingham et al., 1987, Theorem 1.5.2), for all $\delta>0$, there exists $N_{\delta}$ such that for all $x>N_{\delta}$,
@@
$\operatorname{\mathbf{E}}\left[\frac{L(\widehat{m}_{n}-\widehat{T}_{n})}{L(n^{1/2})}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right]$
$\qquad\leq\operatorname{\mathbf{E}}\left[\left(\frac{\widehat{m}_{n}-\widehat{T}_{n}}{n^{1/2}}\right)^{\alpha}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right]$
$\qquad\qquad+\delta\operatorname{\mathbf{E}}\left[\left(\frac{\widehat{m}_{n}-\widehat{T}_{n}}{n^{1/2}}\right)^{\alpha-\rho}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right].$
@@
Using e.g. (Aïdékon and Shi, 2010, Lemma 4.1), there exist $C>0$ and $h>0$ such that for all $n\in\mathbb{N}$, $a\geq 0$ and $b\geq-a$, we have
$\operatorname{\mathbf{P}}(-\widehat{T}_{n}\in[b,b+h],-\widehat{T}_{j}\geq-a,j\leq n)\leq C\frac{\left((a+1)\wedge n^{1/2}\right)\left((a+b+1)\wedge n^{1/2}\right)}{n^{3/2}}.$
@@
$\operatorname{\mathbf{E}}\left[\left(\widehat{m}_{n}-\widehat{T}_{n}\right)^{\gamma}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right]$
$\qquad\leq\sum_{k=1}^{{\left\lceil\varepsilon n^{1/2}/h\right\rceil}}(kh)^{\gamma}\operatorname{\mathbf{P}}(\widehat{m}_{n}-\widehat{T}_{n}\in[kh,(k+1)h],-\widehat{T}_{j}\geq-A,j\leq n)$
$\qquad\leq\frac{Ch^{\gamma}(A+1)}{n^{3/2}}\sum_{k=1}^{{\left\lceil\varepsilon n^{1/2}/h\right\rceil}}k^{\gamma}(A+kh+C^{\prime}\log n+1)$
$\qquad\leq\frac{2Ch^{\gamma}(A+1)^{2}}{n^{3/2}}\sum_{k=1}^{{\left\lceil\varepsilon n^{1/2}/h\right\rceil}}k^{\gamma}(kh+C^{\prime}\log n),$